OpenStudy (anonymous):

Point A is located at (-2, -6), and D is located at (-6, 8). Find the coordinates of the point that lies halfway between A and D. (5 points) Select one: a. (0, 2) b. (-2, 3) c. (-3, 0) d. (-4, 1)

OpenStudy (anonymous):

@ganeshie8 @iambatman

OpenStudy (anonymous):

\[\huge Midpoint =\left( \frac{ x _{1}+x _{2} }{ 2 },\frac{ y _{1}+y _{2} }{ 2 }\right)\]

OpenStudy (anonymous):

So just plug in the values, and you're done.

OpenStudy (anonymous):

ok thnx

OpenStudy (anonymous):

ok let me do it first then check ok

OpenStudy (anonymous):

-4,1

OpenStudy (anonymous):

Yeah man